OpenStudy (anonymous):

the answer of this function is :A+B+C after simplification .i have tried to solve it but failed .is there anyone know how to solve this boolean function

OpenStudy (radar):

This represents a three input "or" circuit.

OpenStudy (radar):

A or B or C What are you trying to do with it? It could also be a more complicated Nand with all three inputs negated (not)

OpenStudy (anonymous):

Did not see you posted a second message, I think this is right (not 100% sure though):\[\overline{a.b+\bar c}+a.\bar c+b = \overline{a.b}.c+a.\bar c+b=(\bar a +\bar b).c+a.\bar c+b=\bar a.c+\bar b.c+a.\bar c +b\]

OpenStudy (anonymous):

haha, oups, I forgot the rest ^^ \[=(\bar a.c+a.\bar c) +(\bar b.c+b)=a \oplus c +b +c = (a \oplus c +c) +b= a + c +b\]